Alyssa wants to adjust the amount of milk and grape juice she drinks each day to control the amount of sodium and carbohydrates in her diet. An ounce of milk contains 15mg of sodium and 1.5g of carbohydrates, while an ounce of grape juice contains 2.5mg of sodium and 5g of carbohydrates. How many ounces of milk and grape juice should Alyssa drink each day to provide 270 mg of sodium and 84g of carbohydrates from these two drinks.
Answer by htmentor(1343)

Let m = the number of ounces of milk
Let j = the number of ounces of juice
The daily intake of sodium will be 15m + 2.5j = 270 [1]
The daily intake of carbohydrates will be 1.5m + 5j = 84 [2]
We have two equations and two unknowns
Multiply [1] by 2 and subtract [2]:
28.5m = 540 - 84 = 456
This gives m = 16
Substitute into [2]:
1.5*16 + 5j = 84
5j = 60
j = 12
So she needs to drink 16 ounces of milk and 12 ounces of juice.